One leg of a right triangle is 4cm shorter thatn the other leg. The length of the hypotenuse is 20 cm. Find the length of the shorter leg.
A. 16cm
B. 17cm
C. 13cm
D. 12cm

Answer by wuwei96815(245)

Right triangle formula is a^2 + b^2 = c^2.

(x)^2 + (x-4)^2 = (20)^2
x^2 + x^2 -8x +16 = 400
2x^2 -8x + 16 = 400
2(x^2-4x+8) = 400
x^2-4x+8 = 200
x^2-4x-192 = 0
(x-16)(x+12) =0

The solutions for x are 16 and -12. We can ignore the -12 answer. So, x=16 and the shorter leg is x-4 or 12.

I hope that you were able to follow the solution.